An initial dose of verapamil is administered to a patient over 2 minutes by slow IV injection. A typical high dose is 0.10 mg per kilogram of body mass, with a maximum dose of 10 mg. For a patient with a weight of 187 pounds, what is the appropriate dose of verapamil?
Answer #1

Answer:

So, 187 pound = 187 / 2.205

=84.80 kilograms.

1kilogram of body mass needs 0.10mg of the drug

Therefore, 84.80 kilograms of body weight will require

= 84.80 × 0.10 = 8.48 mg.

The appropriate dose of verapamil is 8.48 mg.